OpenStudy (djbreezy):

What is the ratio of x to y 5X=6y A.) 6/5 B.) 5/6 C.) 1/5 D.) 1/6

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (djbreezy):

Its A....

OpenStudy (djbreezy):

Ik its A

OpenStudy (djbreezy):

I'm 100% sure

OpenStudy (skullpatrol):

$$5x=6y$$ $$\text{want: } \dfrac{x}{y}$$

OpenStudy (djbreezy):

The test says A is correct

OpenStudy (skullpatrol):

Prove it.

OpenStudy (skullpatrol):

$$\huge\dfrac{1}{5y}•5x=\dfrac{1}{5y}•6y$$